Graph iz(t) for t o tre M 4 65 Helllle 래 24h 2 1 12e А
Answer #1

At taot in transient state, Enductor act as open. Circut 12 qur ilot) = oA At tanya in steady sease, Enductor acts as short cTo find ? or mi 122 Ray heq heq = 6+ ( 4X12 ] = 6+3 = 95 L 4+12) pe thing